If the average of 4 tests is 76, and you take away the highest score, leaving the average of 3 tests 72, what is the highest score

Found 2 solutions by richwmiller, greenestamps:
Answer by richwmiller(17219)

(a+b+c+d)/4=76
(a+b+c+d)=4*76=304
(a+b+c)/3=72
(a+b+c)=3*72=216
d=304-216=88 highest score

Answer by greenestamps(13215)


Nearly everybody will solve this kind of problem exactly the way shown by the other tutor.

But here is an alternative method; try it and see if it "works" for you....

The average dropped by 4 points, and now there are only 3 tests.
Compared to the previous average, the total dropped by 4*3=12 points.
That means the high score that was taken away was 12 points above the previous average.
So the high score that was removed was 76+12 = 88.